Make New Entry,
Make Followup Entry
User name R. Michaels
Log entry time 00:59:35 on March31,2012
Entry number 368482
This entry is a followup to: 368467
keyword=various target crises
First, regarding HALOG 368467, the target pol. was zero simply
because we moved to C12. The frequency readback says 140 MHz,
but this is a false reading, according to James Maxwell. It is
more like 70 MHz. When we put the target back in (later) it should
be polarized and we are supposed to spend the night tweaking the
microwaves to maximize the polarization, while ignoring the
false reading of the frequency. I don't really understand it, but
it seems to be a software problem (?) which the experts will fix
in the morning.
Next crisis: The OVS vacuum is going bad, says Chris Keith, who
called in. It is in the 10^-5 range. Possibilities are that the
pump is not running or a faulty ADC readback. He asked me to
notify James Maxwell. I called James, he's looking.
Then at 0:58 Chris called again and said he's coming in.
In the meantime there were 2 or 3 yellow or red alarms about
the refridgerator. I hardly have time to type this.